A force of 500 N acts horizontally on a load at rest on the floor. The load moves a distance of 20 m in 10 s. What is the amount of power?


1
Expert's answer
2022-02-11T08:48:26-0500

Given:

"F=500\:\rm N"

"d=20\:\rm m"

"t=10\:\rm s"


The power

"P=\frac{W}{t}=\frac{Fd}{t}"

"P=\rm\frac{500\: N*20\: m}{10\: s}=1000\:\rm W"